Circular No. 6084

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

PERIODIC COMET LONGMORE (1994q) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery of

this comet as shown below. The comet is essentially of stellar appearance and its position in very close agreement with the prediction on MPC 20123:

1994 UT  α_{2000}  δ_{2000}  m_{1}
Sept.27.47845  9^{h}39^{m}46.^{s}16  +35^{o}27′49′′.7  20.5
27.48339  9 39 46.66  +35 27 47.5
27.48787  9 39 47.15  +35 27 47.4
28.46961  9 41 16.32  +35 23 25.8
28.47579  9 41 16.70  +35 23 24.8
28.48108  9 41 17.30  +35 23 22.8  21.0

X-RAY NOVA IN SCORPIUS B. A. Harmon, C. A. Wilson, W. S. Paciesas, S. N. Zhang and G. J.

Fishman report for the BATSE/Compton Observatory Team: “We have been observing this object in the 20–100-keV band at a level of 75 mCrab (±30 percent) since Sept. 17. This followed a more intense period, which began on Sept. 6 (IAUC 6075) and ended approximately on Sept. 16, when the source showed daily variations between 200 and 900 mCrab. The spectrum continues to be well represented by a power law between 20 and 200 keV with an index of ∼ −3 in the Sept. 6–16 time period.”

GX 339−4

Harmon et al. add: “We are currently observing a hard (low) state outburst of the black-hole candidate GX 339−4. During Sept. 24–26 its intensity in the 20–100-keV band gradually increased from 150 to 200 mCrab (±15 percent). Owing to the recent activity of several hard x-ray sources in the Sco-Ara region, it is difficult at this time to establish when the outburst began. Preliminary indications are that the source was first detectable around Sept. 5–10. Based on the behavior of previous outbursts, the measured thermal bremsstrahlung temperature of kT = 73 ± 8 keV is consistent with the source’s being a few weeks to a month into its outburst.”

NOTICE TO SUBSCRIBERS Subscribers should note that the new e-mail address and FAX number

at the top of this Circular can be used for information about subscriptions but should not be used for scientific matters.

1994 October 1 Brian G. Marsden